the price elasticity of demand for commodity x is _1.2. Therefore if the price of x increases by 50% the quantity demand decrease by the same propotion true or false?
1
Expert's answer
2020-08-28T10:50:28-0400

False


Price elasticity of demand ="\frac{\%change in quantity}{\% change price}" = 1.2


% "\Delta" quantity = % "\Delta" price "\times" 1.2

= 50% "\times 1.2= 60%" %


quantity demanded will decrease by 60 percent and not 50 percent